OpenStudy (anonymous):

Evaluate the definite integral: (X+1)^2 from x=0 to x=-8 I tried using F(b)-F(a) but it didn't work

OpenStudy (anonymous):

I think an easy way to see it is by u substitution. Integrate then evaluate.

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

So if you set u=x+1 then du=dx so your integral: u^(2)du. Integrate that like normal, then when you're done get the answer back in terms of x and use your limits. Or you can change your limits using u=x+1.

OpenStudy (anonymous):

ahhh kk I'll try that

OpenStudy (anonymous):

totally worked thanks man!

OpenStudy (anonymous):

\[\int (1+x)^2 dx =\frac{1}{3} (x+1)^3=F(x)\\ F(-8)-F(0)=-\frac{344}{3} \]

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

can I use that for integrals that have like x(x^2+5)^5?

OpenStudy (anonymous):

its not a definite one but

OpenStudy (anonymous):

nvm we didn't learn those types yet

OpenStudy (anonymous):

getting ahead of myself

OpenStudy (turingtest):

you can use u-substitution for the last integral you posted u=x^2+5 if you still don't understand pleas post this as a separate question thanks :)

OpenStudy (anonymous):

Yes, and when you take the derivative to get du you'll find that things will cancel.